Write an equation of the line parallel to the line given by -4x + 8y = 5 and having the same y-intercept as the line given by 4x - 3y = 0.
Answer by psbhowmick(878)

The line 4x - 3y = 0 passes through the origin so its y-intercept is 0.

To find slope of the line -4x + 8y = 5 we need to express y as a function of x.
-4x + 8y = 5
8y = 4x + 5
y = 0.5x + 0.625

So the slope is 0.5.

Equation to the line with slope 0.5 and having 0 y-intercept is
y = 0.5x
2y = x